# A Mutual Information Perspective on Multiple Latent Variable Generative Models for Positive View Generation

### v1

@@ Abstract
In image generation, Multiple Latent Variable Generative Models (MLVGMs) employ multiple latent variables to gradually shape the final images, from global characteristics to finer and local details (e.g., StyleGAN, NVAE), emerging as powerful tools for diverse applications. Yet their generative dynamics and latent variable utilization remain only empirically observed. In this work, we propose a novel framework to systematically quantify the impact of each latent variable in MLVGMs, using Mutual Information (MI) as a guiding metric. Our analysis reveals underutilized variables and can guide the use of MLVGMs in downstream applications.
With this foundation, we introduce a method for generating synthetic data for Self-Supervised Contrastive Representation Learning (SSCRL). By leveraging the hierarchical and disentangled variables of MLVGMs, and guided by the previous analysis, we apply tailored latent perturbations to produce diverse views for SSCRL, without relying on real data altogether.
Additionally, we introduce a Continuous Sampling (CS) strategy, where the generator dynamically creates new samples during SSCRL training, greatly increasing data variability. Our comprehensive experiments demonstrate the effectiveness of these contributions, showing that MLVGMs’ generated views compete on par with or even surpass views generated from real data.
This work establishes a principled approach to understanding and exploiting MLVGMs, advancing both generative modeling and self-supervised learning.
@@
1 Introduction
Latent Variable Generative Models (LVGMs), including Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), are foundational approaches for image generation. Given a random variable $\mathbf{X}\in\mathcal{X}$, representing high-dimensional pictures in pixel space, LVGMs aim to approximate the underlying data distribution $p(\mathbf{X})$. To achieve this, they learn a parameterized generator $g(\mathbf{z};\theta)=\mathbf{x}$, where $\mathbf{Z}\in\mathcal{Z}$ denotes a latent variable sampled from a simpler and known distribution in the $\mathcal{Z}$ latent space. A key objective of the learning process is to ensure that the generator is continuous, such that neighboring latent points $\mathbf{z}^{\prime}$ and $\mathbf{z}^{\prime\prime}$ are mapped to perceptually similar outputs $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime\prime}$. This regularization of the latent space allows LVGMs to generate novel content and meaningfully interpolate latent features (Radford et al., 2016; Higgins et al., 2017).
Over the years, significant advancements have been made in latent generative modeling architectures (Vahdat & Kautz, 2020; Karras et al., 2019; 2020; 2021; Sauer et al., 2022), particularly through the introduction of multiple latent variables that are progressively integrated into the network (see Figure 1(a)). This hierarchical design enhances control over the generation process, as different latent variables influence distinct aspects of the output image, ranging from coarse-grained global characteristics to finer, localized features (Figure 1(b)). Consequently, modern Multiple Latent Variable Generative Models (MLVGMs) can generate high-resolution images with improved precision and richer detail.
One major benefit of this seemingly simple improvement is the potential application of MLVGMs across various tasks. Notably, the StyleGAN architecture (Karras et al., 2019) has demonstrated exceptional performance in image editing (Alaluf et al., 2022; Pehlivan et al., 2023), manipulation (Tov et al., 2021), and translation (Richardson et al., 2021). Additionally, recent studies have shown that MLVGMs can serve as effective foundation models for tasks such as adversarial purification (Serez et al., 2024). Collectively, these findings highlight the versatility of MLVGMs, showcasing their utility not only in creative and generative domains but also as pre-trained models for broader applications.
Nevertheless, existing research primarily leverages the “global-to-local” behavior of MLVGMs as an empirical tool, applying it across diverse tasks without delving into the mechanics of latent variable utilization. Despite their effectiveness, these methods fail to investigate how each latent variable, introduced at different stages of the generative process, contributes to the hierarchical refinement of global and local image features. To address this gap, we propose a novel approach that establishes a direct relationship between feature distances in each latent space ($\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,\mathbf{Z}_{n}$) and mutual information (MI) shifts in the shared image space $\mathbf{X}$. Our analysis reveals that achieving equivalent MI shifts by varying a single latent variable (e.g. $\mathbf{Z}_{i}$ to $\hat{\mathbf{Z}}_{i}$) requires progressively larger average perturbations ($\mu_{i}$ see section 3.1) as the $i$-th variable is introduced later in the generative process. This observation aligns with the intuition that the influence of individual latent variables diminishes throughout the generative hierarchy (see Figure 2).
Our systematic evaluation offers a quantitative framework for understanding the contribution of each latent variable in MLVGMs, extending beyond empirical observations. Notably, we also find that the latter latent variables are often underutilized in modern MLVGMs, highlighting potential inefficiencies in current training paradigms. With these insights, we not only provide a deeper understanding of MLVGMs but also can guide their effective utilization in downstream tasks.
@@ 1 Introduction
Building on this understanding, we propose a novel application of MLVGMs in Self-Supervised Contrastive Representation Learning (SSCRL). In SSCRL, feature extractors, or encoders $f(\mathbf{x};\phi)$ with parameters $\phi$, are trained to represent data by contrasting positive and negative views. Positive views are semantically similar images, encouraged to have close representations in the latent space, while negative views correspond to unrelated data points that are forced to have distant representations. Therefore, we propose to leverage the different impacts of multiple latent variables in MLVGMs to manipulate specific features and generate positive views. This approach enables the training of SSCRL encoders without relying on real data, demonstrating the potential of MLVGMs as pre-trained models for producing high-quality synthetic images tailored for representation learning.
@@
The primary objective of SSCRL is to enforce a desired set of invariances in the learned representations (Xiao et al., 2020), achieved by creating valid positive views. Figure 3 compares the proposed method with standard pixel-space augmentations and single latent variable generative models (LVGMs) for view generation. In the typical approach (Figure 3(a)), a finite set of hand-crafted transformations, such as color adjustments, cropping, or flipping, is applied directly in the pixel space. Alternatively, invariances can be introduced at the latent level of a pre-trained LVGM (Figure 3(b)). However, in LVGMs, all image features are entangled within a single latent space, making it difficult to generate specific invariances (e.g., altering fur patterns) without inadvertently affecting global features, such as changing the dog breed (e.g., from Australian Terrier to Yorkshire Terrier in the figure). In contrast, MLVGMs inherently disentangle global and local features, enabling precise control over specific characteristics. For instance, using MLVGMs, attributes like fur patterns or color can be modified while preserving global features, such as the dog breed. This is achieved by independently perturbing each latent variable to a desired magnitude, as illustrated in Figure 3(c).
@@
To evaluate our contributions, we apply our novel quantification algorithm to two distinct MLVGMs: a BigBiGAN (Donahue & Simonyan, 2019) pre-trained on ImageNet-$1$K (Deng et al., 2009), and a StyleGAN2 (Karras et al., 2020) pre-trained on LSUN Cars (Yu et al., 2015). Subsequently, we leverage the same MLVGMs to generate views for SSCRL using our proposed Continuous Sampling (CS) strategy. Specifically, we train feature extractors with three different SSCRL framweorks, namley SimCLR (Chen et al., 2020), SimSiam (Chen & He, 2021), and BYOL (Grill et al., 2020). The learned representations are then validated following a consolidated practice through linear classification across multiple downstream datasets and object detection on Pascal VOC (Everingham et al., 2010). Our results demonstrate that MLVGM-based view generation outperforms simple LVGM-based techniques and achieves comparable or superior results than training with real data. Additionally, we measure the training time per epoch using Continuous Sampling vs. standard data loading, establishing CS as an efficient alternative for increasing data diversity.
@@ 2 Related Works
The idea of utilizing multiple latent variables is well-established in the generative models’ literature, typically presented as an evolution of Latent Variable Generative Models (LVGMs). For instance, Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) leverage multiple latent variables to enhance the expressivity of approximate distributions, as demonstrated by architectures such as NVAE (Vahdat & Kautz, 2020) and Ladder VAE (Child, 2020), or to improve latent disentanglement, as in Li et al. (2019). Similarly, Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have embraced this concept in models like LapGAN (Denton et al., 2015), BigGANs (Brock et al., 2018; Donahue & Simonyan, 2019), and GigaGAN (Kang et al., 2023). Advances in Normalizing Flows (Dinh et al., 2015; Rezende & Mohamed, 2015) have also incorporated multiple latent variables, with works like Hu et al. (2022) introducing architectures inspired by physics to achieve this goal.
In such a rich environment, numerous applications of MLVGMs have emerged. Of particular relevance in this context is the StyleGAN family (Karras et al., 2019; 2020; 2021; Sauer et al., 2022), which has been widely applied in image editing and manipulation tasks (Tov et al., 2021; Richardson et al., 2021; Alaluf et al., 2022; Pehlivan et al., 2023). More recently, MLVGMs have also been used as foundation models for non-generative downstream tasks, such as purification against adversarial attacks (Serez et al., 2024).
Motivated by this growing body of research, in this work, we address the problem of quantitatively evaluating the impact of individual latent variables in MLVGMs. To this end, we propose a novel method based on information theory, capable of revealing underutilized variables and serving as a tool for further applications. Furthermore, we employ MLVGMs in an unprecedented way, that is to generate positive views for Self-Supervised Contrastive Representation Learning (SSCRL).
@@
Self-Supervised Contrastive Representation Learning (SSCRL) (Hadsell et al., 2006) aims to learn meaningful latent representations without relying on labeled data, primarily by designing informative positive views (Tian et al., 2020; Xiao et al., 2020). Early approaches, such as Bachman et al. (2019); Misra & Maaten (2020); Caron et al. (2020), focused on pretext tasks like matching global and local parts of an image to create multiple views. Subsequently, SimCLR (Chen et al., 2020), a foundational method in the field, introduced the use of manually designed transformations, including flipping, cropping, and color distortions. More recent works have explored advanced techniques, such as learning views in an adversarial manner (Tamkin et al., 2020; Shi et al., 2022) or projecting anchor images into the latent spaces of pre-trained generators (Yang et al., 2022; Astolfi et al., 2023; Kim et al., 2023; Wu et al., 2023; Han et al., 2023; Zeng et al., 2024).
The unifying feature of these methods is their reliance on real datasets, using anchor images as the starting point. In contrast, our approach generates views exclusively from synthetic data using MLVGMs. As a consequence, our method is complementary to existing techniques, as transformations like those introduced by SimCLR (denoted as $T_{\mathbf{x}}(\mathbf{x})$ in Figure 3(a)) can be seamlessly applied to the views generated by MLVGMs, potentially leading to even more informative representations. We test this hypothesis in the experimental section, combining pixel-space augmentations with our latent space views.
Further along our line of work, methods like Jahanian et al. (2021); Li et al. (2022) have proposed generating fully synthetic views by sampling nearby points in the latent space of pre-trained LVGMs (Figure 3(b)). However, the primary limitation of these methods lies in the entanglement of all image features within a single latent space, which complicates the task of generating valid positive views. By leveraging the multiple latent spaces of MLVGMs, our approach disentangles coarse, global information from finer, local details, greatly simplifying the definition of valid views and improving the quality of the learned representations (Figure 3(c)).
@@
The primary challenge in these approaches is the limited diversity of generated data, which has been identified as a key factor contributing to the performance gap between classifiers trained on real versus synthetic datasets (Ravuri & Vinyals, 2019). Recent studies (Fan et al., 2024) suggest that scaling up the size of synthetic training sets can reduce this accuracy gap, though it does not fully eliminate it. However, generating large datasets introduces its own set of challenges, particularly increased disk space usage and data management overhead. Existing methods (Besnier et al., 2020; Lampis et al., 2023) address this issue by partially renewing synthetic data at each epoch or by regenerating the dataset entirely every $N$ epochs.
In contrast, we leverage fast-sampling models, such as GANs, to generate data directly during training. This method, referred to as Continuous Sampling (CS), eliminates the need for disk storage, avoids the bottleneck of data loading, and ensures competitive training times. More importantly, CS provides a continuous stream of fresh images at every training step, maximizing data diversity and effectively addressing the limitations of prior techniques.
@@ 3.1 Measuring the impact of latent variables in MLVGMs
The recent success of Multiple Latent Variable Generative Models (MLVGMs) in diverse applications (Tov et al., 2021; Richardson et al., 2021; Alaluf et al., 2022; Pehlivan et al., 2023; Serez et al., 2024) underscores the need for a systematic method to quantify the contribution of each latent variable in the generative process. Developing such a method would enhance our understanding of the hierarchical dynamics of MLVGMs, identifying underutilized or overutilized latent codes and offering valuable insights for optimizing their application in downstream tasks. To formalize our approach, we first define the concept of MLVGMs:
@@
Each latent variable $\mathbf{z}_{i}$ contributes differently to the generation of the final output $\mathbf{x}$, depending on its role in the generative process. To compare these contributions meaningfully, we require a metric that operates in a common space. Since the output $\mathbf{x}$ resides in the high-dimensional pixel space and represents a random variable $\mathbf{X}$, we select Mutual Information (MI) as the metric of choice [^1] See Appendix A for the formal definition of Mutual Information and its probabilistic interpretation..
Intuitively, consider an MLVGM with $n=3$ latent variables, as shown in Figure 1. Let $\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}$ represent the random latent variables for the three latent spaces, and $\mathbf{X}$ the output in the pixel space. Suppose we perturb the first latent variable, replacing $\mathbf{Z}_{1}$ with $\hat{\mathbf{Z}}_{1}$. This generates a modified random variable $\mathbf{X}^{\prime}$ in the pixel space. We can now relate the average magnitude of the perturbation in the latent space (e.g. using $\text{L}_{2}$ distance), $\mu_{1}=\mathbb{E}[\|\hat{\mathbf{Z}}_{1}-\mathbf{Z}_{1}\|_{2}]$, to the resulting Mutual Information shift in the pixel space, $I(\mathbf{X},\mathbf{X}^{\prime})=\gamma$.
The same process can be repeated for $\mathbf{Z}_{2}$ and $\mathbf{Z}_{3}$, introducing $\hat{\mathbf{Z}}_{2}$ and $\hat{\mathbf{Z}}_{3}$, and calculating the perturbation magnitudes $\mu_{2}$ and $\mu_{3}$ needed to achieve the same MI shift $\gamma$ in the pixel space. If the generative process respects the “global-to-local” hierarchy typically attributed to MLVGMs (Figure 1(b)), we expect: $\mu_{3}>\mu_{2}>\mu_{1}$, as depicted in Figure 2.
@@
In typical SSCRL setups (e.g., SimCLR (Chen et al., 2020)), the random variables $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are generated using deterministic augmentations, such as cropping, flipping, or color adjustment, applied to the same base image. These transformations result in a fixed mutual information value $I(\mathbf{X};\mathbf{X}^{\prime})$. Since $f(\cdot;\phi)$ is a deterministic function, the fixed term $I(\mathbf{X};\mathbf{X}^{\prime})$ serves as an upper bound to Equation 2, following directly from the data processing inequality (see Appendix A):
@@
Estimating MI in MLVGMs.
The proposed algorithm builds on the insights of Equation 3, utilizing InfoNCE as a proxy to measure MI shifts between views. Unlike classical SSCRL methods, which rely on fixed, deterministic transformations, we generate views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by perturbing individual latent variables in the latent spaces of a pre-trained MLVGM.
@@
For the training procedure, we initialize $\varphi$ such that $T_{\mathbf{z}}^{i}(\cdot)$ represents the identity function. In this state, the perturbation is zero, and the views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are identical. As a result, from the perspective of Equation 3, the mutual information between the views is at its maximum, $I(\mathbf{X};\mathbf{X}^{\prime})=H(\mathbf{X})$, which corresponds to a trivial setting where the encoder $f$ can achieve $\mathcal{L}_{\text{InfoNCE}}\approx 0$ with ease. As training progresses, the perturbation function $T_{\mathbf{z}}^{i}(\cdot)$ learns to apply progressively larger modifications to the latent variable $\mathbf{Z}_{i}$, increasing the diversity of the generated views. This, in turn, reduces the mutual information $I(\mathbf{X};\mathbf{X}^{\prime})$, thereby lowering the upper bound in Equation 3. As a result, the encoder $f$, tasked with minimizing $\mathcal{L}_{\text{InfoNCE}}$, must maintain the shared information between increasingly distinct views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ into a common representation, tightening the lower bound.
In summary, $T_{\mathbf{z}}^{i}$ progressively enhances diversity in the views, reducing $I(\mathbf{X};\mathbf{X}^{\prime})$ and causing InfoNCE to increase over time. Conversely, $f$ seeks to learn the most informative representations, tightening the lower bound from the left and seeking equality in Equation 3. We refer the reader to Appendix B for a detailed graphical illustration of these training dynamics, showing the evolution of InfoNCE loss, average perturbations, and additional insights such as required training time and hyperparameter settings.

### v2

@@ Abstract
In image generation, Multiple Latent Variable Generative Models (MLVGMs) employ multiple latent variables to gradually shape the final images, from global characteristics to finer and local details (e.g., StyleGAN, NVAE), emerging as powerful tools for diverse applications. Yet their generative dynamics remain only empirically observed, without a systematic understanding of each latent variable’s impact. In this work, we propose a novel framework that quantifies the contribution of each latent variable using Mutual Information (MI) as a metric. Our analysis reveals that current MLVGMs often underutilize some latent variables, and provides actionable insights for their use in downstream applications.
With this foundation, we introduce a method for generating synthetic data for Self-Supervised Contrastive Representation Learning (SSCRL). By leveraging the hierarchical and disentangled variables of MLVGMs, our approach produces diverse and semantically meaningful views without the need for real image data. Additionally, we introduce a Continuous Sampling (CS) strategy, where the generator dynamically creates new samples during SSCRL training, greatly increasing data variability. Our comprehensive experiments demonstrate the effectiveness of these contributions, showing that MLVGMs’ generated views compete on par with or even surpass views generated from real data.
This work establishes a principled approach to understanding and exploiting MLVGMs, advancing both generative modeling and self-supervised learning. Code and pre-trained models at: https://github.com/SerezD/mi_ml_gen.
@@
1 Introduction
Latent Variable Generative Models (LVGMs), such as Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), are foundational approaches for image generation. Given a random variable $\mathbf{X}\in\mathcal{X}$, representing high-dimensional pictures in pixel space, LVGMs aim to approximate the underlying data distribution $p(\mathbf{X})$. To achieve this, they learn a parameterized generator $g(\mathbf{z};\theta)=\mathbf{x}$, where $\mathbf{Z}\in\mathcal{Z}$ denotes a latent variable sampled from a simpler and known distribution in the $\mathcal{Z}$ latent space. A key objective of the learning process is to ensure that the generator is continuous, such that neighboring latent points $\mathbf{z}^{\prime}$ and $\mathbf{z}^{\prime\prime}$ are mapped to perceptually similar outputs $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime\prime}$. This regularization of the latent space allows LVGMs to generate novel content and meaningfully interpolate latent features (Radford et al., 2016; Higgins et al., 2017).
Over the years, advancements in latent generative modeling have focused on the use of multiple latent variables, rather than a single latent code (Vahdat & Kautz, 2020; Karras et al., 2019; 2020; 2021; Sauer et al., 2022). By incorporating latent variables at different layers of the network (see Figure˜1(a)), these Multiple Latent Variable Generative Models (MLVGMs) offer a hierarchical structure where early latent codes influence broad, global features and later codes refine finer, local details (Figure˜1(b)). The resulting architecture enhances control over the image synthesis process, enabling the generation of high-resolution images with richer detail and improved precision.
The ability to disentangle global and local features in image generation not only improves the visual quality of generated images but also widens the application scope of these models. For instance, the StyleGAN architecture (Karras et al., 2019) has demonstrated exceptional performance in image editing (Alaluf et al., 2022; Pehlivan et al., 2023), manipulation (Tov et al., 2021), and translation (Richardson et al., 2021). Additionally, recent studies have shown that MLVGMs can serve as effective foundation models for tasks such as adversarial purification (Serez et al., 2025). Collectively, these findings highlight the versatility of MLVGMs, showcasing their utility not only in creative and generative domains but also as pre-trained models for broader applications.
Nevertheless, existing research primarily leverages the “global-to-local” behavior of MLVGMs as an empirical tool, applying it across diverse tasks without delving into the mechanics of latent variable utilization. In other terms, these approaches assume that earlier latent variables shape coarse image attributes while later ones refine fine details, but they do so without formally analyzing how each latent variable contributes to image generation. As a result, the internal dynamics of MLVGMs remain poorly understood.
To address this gap, we propose a novel approach that establishes a direct relationship between feature distances in each latent space ($\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,\mathbf{Z}_{n}$) and mutual information (MI) shifts in the shared image space $\mathbf{X}$. The key insight is that producing an equivalent MI shift in the output space requires increasingly larger perturbations ($\mu_{i}$) as we move deeper into the generative hierarchy—i.e., from $\mathbf{Z}_{1}$ to $\mathbf{Z}_{n}$. This quantitatively confirms the global-to-local pattern and reveals how influence diminishes across successive latent variables (see Figure˜2).
@@ 1 Introduction
By grounding this analysis in information theory, our approach moves beyond intuition and provides a principled framework for understanding latent variable roles in MLVGMs. It enables direct comparisons between latent spaces and exposes inefficiencies in how models allocate representational power. Notably, we reveal the underutilization of later-stage latents in all tested models, opening to further architectural improvements and enabling better strategies for leveraging MLVGMs in downstream tasks.
@@
With this understanding, we propose a novel application of MLVGMs in Self-Supervised Contrastive Representation Learning (SSCRL). In SSCRL, feature extractors, or encoders $f(\mathbf{x};\phi)$ with parameters $\phi$, are trained to represent data by contrasting positive and negative views. Positive views are semantically similar images, encouraged to have close representations in the latent space, while negative views correspond to unrelated data points that are forced to have distant representations. Therefore, we propose to leverage the different impacts of multiple latent variables in MLVGMs to manipulate specific features and generate positive views. This approach enables the training of SSCRL encoders without relying on real data, demonstrating the potential of MLVGMs as pre-trained models for producing high-quality synthetic images tailored for representation learning.
@@
The primary objective of SSCRL is to enforce a desired set of invariances in the learned representations (Xiao et al., 2020), achieved by creating valid positive views. Figure˜3 compares the proposed method with standard pixel-space augmentations and single latent variable generative models (LVGMs) for view generation. In the typical approach (Figure˜3(a)), a finite set of hand-crafted transformations, such as color adjustments, cropping, or flipping, is applied directly in the pixel space. Alternatively, invariances can be introduced at the latent level of a pre-trained LVGM (Figure˜3(b)). However, in LVGMs, all image features are entangled within a single latent space, making it difficult to generate specific invariances (e.g., altering fur patterns) without inadvertently affecting global features, such as changing the dog breed (e.g., from Australian Terrier to Yorkshire Terrier in the figure). In contrast, MLVGMs inherently disentangle global and local features, enabling precise control over specific characteristics. For instance, using MLVGMs, attributes like fur patterns or color can be modified while preserving global features, such as the dog breed. This is achieved by independently perturbing each latent variable to a desired magnitude, as illustrated in Figure˜3(c).
@@
To evaluate our contributions, we apply our novel quantification algorithm to two distinct MLVGMs: a BigBiGAN (Donahue & Simonyan, 2019) pre-trained on ImageNet-11K (Deng et al., 2009), and a StyleGAN2 (Karras et al., 2020) pre-trained on LSUN Cars (Yu et al., 2015). Subsequently, we leverage the same MLVGMs to generate views for SSCRL using our proposed Continuous Sampling (CS) strategy. Specifically, we train feature extractors with three different SSCRL frameworks, namely SimCLR (Chen et al., 2020), SimSiam (Chen & He, 2021), and BYOL (Grill et al., 2020). The learned representations are then validated following a consolidated practice through linear classification across multiple downstream datasets and object detection on Pascal VOC (Everingham et al., 2010). Our results demonstrate that MLVGM-based view generation outperforms simple LVGM-based techniques and achieves comparable or superior results than training with real data. Additionally, we measure the training time per epoch using Continuous Sampling vs. standard data loading, establishing CS as an efficient alternative for increasing data diversity.
@@ 2 Related Works
The idea of utilizing multiple latent variables is well-established in the generative models’ literature, typically presented as an evolution of Latent Variable Generative Models (LVGMs). For instance, Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) leverage multiple latent variables to enhance the expressivity of approximate distributions, as demonstrated by architectures such as NVAE (Vahdat & Kautz, 2020) and VD-VAE (Child, 2021), or to improve latent disentanglement, as in Li et al. (2019). Similarly, Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have embraced this concept in models like LapGAN (Denton et al., 2015), BigGANs (Brock et al., 2018; Donahue & Simonyan, 2019), and GigaGAN (Kang et al., 2023). Advances in Normalizing Flows (Dinh et al., 2015; Rezende & Mohamed, 2015) have also incorporated multiple latent variables, with works like Hu et al. (2022) introducing architectures inspired by physics to achieve this goal.
In such a growing environment, numerous applications of MLVGMs have emerged. Of particular relevance in this context is the StyleGAN family (Karras et al., 2019; 2020; 2021; Sauer et al., 2022), which has been widely applied in image editing and manipulation tasks (Tov et al., 2021; Richardson et al., 2021; Alaluf et al., 2022; Pehlivan et al., 2023). More recently, MLVGMs have also been used as foundation models for non-generative downstream tasks, such as purification against adversarial attacks (Serez et al., 2025). Motivated by such a rapid expansion, in this work we propose a novel application in SSCRL and for the first time study how the “generative load” is spread across latent variables in the current architectures.
@@
Self-Supervised Contrastive Representation Learning (SSCRL) (Hadsell et al., 2006) aims to learn meaningful latent representations without relying on labeled data, primarily by designing informative positive views (Tian et al., 2020; Xiao et al., 2020). Early approaches, such as Bachman et al. (2019); Misra & Maaten (2020); Caron et al. (2020), focused on pretext tasks like matching global and local parts of an image to create multiple views. Subsequently, SimCLR (Chen et al., 2020), a foundational method in the field, introduced the use of manually designed transformations, including flipping, cropping, and color distortions. More recent works have explored advanced techniques, such as learning views in an adversarial manner (Tamkin et al., 2020; Shi et al., 2022) or projecting anchor images into the latent spaces of pre-trained generators (Yang et al., 2022; Astolfi et al., 2023; Kim et al., 2023; Wu et al., 2023; Han et al., 2023; Zeng et al., 2024). While our approach differs by relying solely on synthetic data, we show that it is also complementary to previous techniques, allowing to apply different transformations (denoted as $T_{\mathbf{x}}(\mathbf{x})$ in Figure˜3(a)) on top of the generated views-i.e. directly in the pixel space.
Further along our line of work, methods like Jahanian et al. (2021); Li et al. (2022) have proposed generating fully synthetic views by sampling nearby points in the latent space of pre-trained LVGMs (Figure˜3(b)). We direct compare against these baselines in the experimental section, showing that the coarse-to-fine feature disentanglement of MLVGMs allows to obtain better views for most downstream tasks.
@@
The primary challenge in these approaches is the limited diversity of generated data, which has been identified as a key factor contributing to the performance gap between classifiers trained on real versus synthetic datasets (Ravuri & Vinyals, 2019). Recent studies (Fan et al., 2024) suggest that scaling up the size of synthetic training sets can reduce this accuracy gap, though it does not fully eliminate it. However, generating large datasets introduces its own set of challenges, particularly increased disk space usage and data management overhead. Existing methods (Besnier et al., 2020; Lampis et al., 2023) address this issue by partially renewing synthetic data at each epoch or by regenerating the dataset entirely every $N$ epochs. In this work, we surpass the common assumption about the inefficiencies of generating training images in real time, leveraging fast-sampling models, such as GANs, to generate data directly during training and avoiding storage and loading bottlenecks altogether.
@@ 3.1 Measuring the impact of latent variables in MLVGMs
Before formalizing our approach for measuring the contribution of single latent variables, we define the concept of Multiple Latent Variable Generative Models (MLVGMs):
@@
Intuitions.
To meaningfully compare the contribution of each latent variable $\mathbf{Z}_{i}$, we need a metric that operates in the common pixel-space. Probabilistically, we consider images as a random variable $\mathbf{X}$, and therefore
select Mutual Information (MI) as the metric of choice [^1] See Appendix A for the formal definition of Mutual Information and its probabilistic interpretation..
As an example, let’s consider an MLVGM with $n=3$ latent variables, as shown in Figure˜1. Let $\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}$ represent the random latent variables for the three latent spaces, and $\mathbf{X}$ the output in the pixel space. Suppose we perturb the first latent variable, replacing $\mathbf{Z}_{1}$ with $\hat{\mathbf{Z}}_{1}$. This generates a modified random variable $\mathbf{X}^{\prime}$ in the pixel space. We can now relate the average magnitude of the perturbation in the latent space (e.g. using $\text{L}_{2}$ distance), $\mu_{1}=\mathbb{E}[\|\hat{\mathbf{Z}}_{1}-\mathbf{Z}_{1}\|_{2}]$, to the resulting Mutual Information shift in the pixel space, $I(\mathbf{X},\mathbf{X}^{\prime})=\gamma$.
The same process can be repeated for $\mathbf{Z}_{2}$ and $\mathbf{Z}_{3}$, introducing $\hat{\mathbf{Z}}_{2}$ and $\hat{\mathbf{Z}}_{3}$, and calculating the perturbation magnitudes $\mu_{2}$ and $\mu_{3}$ needed to achieve the same MI shift $\gamma$ in the pixel space. If the generative process respects the “global-to-local” hierarchy typically attributed to MLVGMs (Figure˜1(b)), we expect: $\mu_{3}>\mu_{2}>\mu_{1}$, as depicted in Figure˜2.
@@
In typical SSCRL setups (e.g., SimCLR (Chen et al., 2020)), the random variables $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are generated using deterministic augmentations, such as cropping, flipping, or color adjustment, applied to the same base image. These transformations result in a fixed mutual information value $I(\mathbf{X};\mathbf{X}^{\prime})$. Since $f(\cdot;\phi)$ is a deterministic function, the fixed term $I(\mathbf{X};\mathbf{X}^{\prime})$ serves as an upper bound to Equation˜2, following directly from the data processing inequality (see Appendix˜A):
@@
The proposed approach.
We build on the insights of Equation˜3, utilizing InfoNCE as a proxy to measure MI shifts between views. Unlike classical SSCRL methods, which rely on fixed, deterministic transformations, we generate views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by perturbing individual latent variables in the latent spaces of a pre-trained MLVGM.
@@
Training dynamics.
We initialize parameters $\varphi$ such that $T_{\mathbf{z}}^{i}(\cdot)$ represents the identity function. In other terms, the applied perturbation is initially zero, and the views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are identical. From the perspective of Equation˜3, $I(\mathbf{X};\mathbf{X}^{\prime})=H(\mathbf{X})$, corresponding to the trivial setting where the encoder $f$ can achieve $\mathcal{L}_{\text{InfoNCE}}\approx 0$ with ease. As training continues, the perturbation function $T_{\mathbf{z}}^{i}(\cdot)$ learns to apply progressively larger modifications to the latent variable $\mathbf{Z}_{i}$, increasing the diversity of the generated views. This, in turn, reduces the mutual information $I(\mathbf{X};\mathbf{X}^{\prime})$, thereby lowering the upper bound in Equation˜3. As a result, the encoder $f$, tasked with minimizing $\mathcal{L}_{\text{InfoNCE}}$, must maintain the shared information between increasingly distinct views $\mathbf{X}$ and $\mathbf{X}^{\prime}$ into a common representation, tightening the lower bound.
In summary, $T_{\mathbf{z}}^{i}$ progressively enhances diversity in the views, reducing $I(\mathbf{X};\mathbf{X}^{\prime})$ and causing InfoNCE to increase over time. Conversely, $f$ seeks to learn the most informative representations, tightening the lower bound from the left and seeking equality in Equation˜3. We refer the reader to Appendix˜B for a detailed graphical illustration of these training dynamics, showing the evolution of InfoNCE loss, average perturbations, and additional insights such as required training time and hyperparameter settings.